# Scaling to Multimodal and Multichannel Heart Sound Classification: Fine-Tuning Wav2Vec 2.0 with Synthetic and Augmented Biosignals

### v2

@@ Abstract
Cardiovascular diseases (CVDs) are the leading cause of death worldwide, accounting for approximately 17.9 million deaths each year. Early detection is critical, creating a demand for accurate and inexpensive pre-screening methods. Deep learning has recently been applied to classify abnormal heart sounds indicative of CVDs using synchronised phonocardiogram (PCG) and electrocardiogram (ECG) signals, as well as multichannel PCG (mPCG). However, state-of-the-art architectures remain underutilised due to the limited availability of synchronised and multichannel datasets. Augmented datasets and pre-trained models provide a pathway to overcome these limitations, enabling transformer-based architectures to be trained effectively. This work combines traditional signal processing with denoising diffusion models, WaveGrad and DiffWave, to create an augmented dataset to fine-tune a Wav2Vec 2.0-based classifier on multimodal and multichannel heart sound datasets. The approach achieves state-of-the-art performance. On the Computing in Cardiology (CinC) 2016 dataset of single channel PCG, accuracy, unweighted average recall (UAR), sensitivity, specificity and Matthew’s correlation coefficient (MCC) reach 92.48%, 93.05%, 93.63%, 92.48%, and 0.8283, respectively. Using the synchronised PCG and ECG signals of the training-a dataset from CinC, 93.14%, 92.21%, 94.35%, 90.10%, and 0.8380 are achieved for accuracy, UAR, sensitivity, specificity and MCC, respectively. Using a wearable vest dataset consisting of mPCG data, the model achieves 77.13% accuracy, 74.25% UAR, 86.47% sensitivity, 62.04% specificity, and 0.5082 MCC. These results demonstrate the effectiveness of transformer-based models for CVD detection when supported by augmented datasets, highlighting their potential to advance multimodal and multichannel heart sound classification.
@@ 1 Introduction
Cardiovascular diseases (CVDs) are the leading cause of death globally, accounting for approximately 17.9 million deaths each year [1]. Since CVD treatment is most effective when the condition is detected early, there is a pressing need for accurate and affordable pre-screening methods. Cardiac auscultation is one such technique: it is inexpensive, non-invasive, and widely used, relying on physicians listening to heart sounds to detect abnormalities indicative of CVD. However, auscultation yields relatively low diagnostic accuracy, partly because heart sounds often lie near the threshold of human hearing [2, 3, 4].
Recent advances in computer-aided methods have demonstrated promise in improving the accuracy of abnormal heart sound classification. Combining phonocardiogram (PCG) and electrocardiogram (ECG) signals within deep learning frameworks has produced favourable results [5]. However, current datasets suffer from limitations, including class imbalance, low signal-to-noise ratio, and limited size, all hindering robust and accurate classification. Current state-of-the-art (SOTA) approaches often use convolutional neural networks (CNNs), many of which leverage pre-trained image-based architectures with spectrogram inputs. Transformer-based models remain underexplored in this domain, with only Vision Transformers (ViTs) evaluated on spectrogram inputs. Therefore, there is significant potential in applying transformers directly to raw audio signals.
Using raw signals as input offers the advantage of preserving phase information and enabling models to learn features that may be difficult to extract from time-frequency representations. However, learning time-frequency features from raw audio alone can be challenging. Modern architectures such as transformers generally require substantial data to outperform existing approaches [6]. Transformer-based models that operate on raw signals show potential for improved performance, but their data requirements must be addressed.
Recent advances in synthetic audio generation using diffusion models offer a promising solution to this data scarcity, enabling easier training of data-hungry models and potentially yielding SOTA performance [7, 8]. Traditional data augmentation techniques have also been applied to mitigate issues with limited and imbalanced datasets [9]. Furthermore, fine-tuning large pre-trained models provides another avenue for overcoming data limitations [5, 10]. Models like Wav2Vec 2.0 (Wav2Vec2) [11] have shown excellent performance in various speech classification tasks [12], making them strong candidates for transfer learning in PCG and ECG classification.
A recent development includes a wearable device equipped with up to seven PCG sensors, enabling the collection of synchronised multichannel PCG data [13]. This advancement makes it feasible to train models utilising multimodal and multichannel data for abnormal heart sound detection.
@@ 1.1 Phonocardiogram and Electrocardiogram Signals
PCG signals consist of multiple sounds created by sudden changes in blood flow within the heart, causing vibrations [14]. The fundamental heart sounds are S1 and S2. S1 occurs due to isovolumetric ventricular contraction at the beginning of systole, and S2 results from the closure of the aortic and pulmonic valves at the beginning of diastole.
While S1 and S2 are the most audible sounds, other sounds can also be heard, such as the third heart sound (S3), fourth heart sound (S4), systolic ejection click, mid-systolic click, opening snap, and heart murmurs, which result from turbulent, fast-flowing blood [15]. These sounds all lie in the low-frequency range: S1 ranges from 10–140Hz with the highest energy around 25–45Hz; S2 spans 10–200Hz with energy concentrated around 55–75Hz; and S3 and S4 fall between 20–70Hz, though they are less audible. Murmurs, which may indicate CVDs, can be found in a wider frequency range from 25Hz to 400Hz [16], with some extending up to 600Hz but with less energy [17].
ECG signals represent the electrical activity of the heart [18]. An ECG signal comprises P waves, QRS complexes, and T waves, with a U wave occasionally present [19]. These components contain diagnostic information useful in identifying CVDs. ECG signals are typically filtered between 0.5Hz and 40Hz to remove baseline wander and unwanted noise [20]. In patients with coronary artery disease (CAD), symptoms such as T-wave inversion, ST-T abnormalities, left ventricular hypertrophy, and premature ventricular contractions have been documented [21].
Combining PCG and ECG signals has yielded better results than using either signal alone [5], as both contain complementary diagnostic features. Figure 1 illustrates these signals in a patient with mitral valve prolapse.
@@
Combining multiple PCG sensors in different auscultation regions can yield better results than single-channel PCG, as each channel offers a different resolution of the region where the murmur originates [22]. Figure 2 shows a CAD patient with multichannel PCG (mPCG) data.
@@
Wav2Vec 2.0 (Wav2Vec2), developed by Meta, is a speech-to-text model with a transformer-based architecture. It consists of three primary components: a feature encoder, transformer encoder and a quantisation module for pre-training [11].
@@
The quantisation module discretises and then encodes the output using a Gumbel-Softmax function during self-supervised training. Training proceeds in two phases: first, self-supervised training on large unlabelled audio datasets (e.g., LibriSpeech [23]) to learn speech representations; second, fine-tuning with 960 hours of labelled speech data across diverse accents and languages [11].
@@ 1.3 Diffusion Models
Diffusion probabilistic models are generative models that convert a simple distribution (e.g., isotropic Gaussian) into a complex data distribution via a Markov chain process [24, 25]. Training is based on optimising the variational lower bound (ELBO), even when the data likelihood is intractable.
These models, applied in audio and image synthesis, relate to denoising score matching and do not require separate encoder or discriminator networks like VAEs [26] or GANs [27]. This advantage avoids issues such as posterior collapse and mode collapse, making diffusion models particularly effective for high-fidelity audio generation by whitening training data latents with a parameter-free noise process.
@@ 1.3.1 WaveGrad
WaveGrad is a diffusion model for conditional audio synthesis [8]. It employs upsampling (UBlocks) and downsampling (DBlocks) blocks, conditioned on mel-spectrogram inputs, along with Feature-wise linear modulation (FiLM) modules. These blocks resemble those in the GAN-TTS model [28]. Figure 4 shows the overall architecture. Module components are shown in Figure 5. The loss function measures the difference between the added noise in the forward diffusion and the predicted noise during denoising.
@@ 1.3.2 DiffWave
DiffWave is a diffusion model for raw audio synthesis with both conditional and unconditional variants [7]. The model uses 1D convolutions and fully connected layers, with its core comprising bi-directional dilated convolutions and residual connections. It is trained using a single ELBO-based objective without auxiliary losses. Conditional generation uses local conditioning signals and global conditioning via discrete labels.
@@ 2.1.1 Multimodal dataset
The classification dataset is sourced from the 2016 PhysioNet Computing in Cardiology Challenge (CinC), comprising five databases (training-a to training-e) [15]. There are 3153 recordings sourced from 764 patients [15], some of which are very noisy. Of these recordings, 665 are abnormal and 2488 are normal, with each recording lasting 5–120 seconds. Database training-a includes synchronised ECG and PCG recordings; out of 409 recordings, 405 contain both signal types (288 abnormal, 117 normal, hence, we conduct analysis on PCG signal only models using the entire dataset, and PCG-ECG hybrid models using only training-a.
Table 1: Datasets from CinC 2016 (adapted from [15])
@@ 2.1.2 Multichannel dataset
Recordings were obtained from subjects using a multichannel wearable vest with seven phonocardiogram sensors recording at different auscultation sites [22]. Ninety-six subjects were diagnosed with coronary artery disease through angiography, and an additional 61 were subjects without the disease. Of the 61, 21 were control subjects below the age of 35, assumed to be normal, as the risk of CAD is significantly higher in people aged 45 and over. The multichannel device had sixty seconds of data recorded in a hospital with background noise, with normal breathing occuring. This vest dataset also did not have optimal positioning of all the stethoscopes, resulting in lower signal-to-noise ratio (SNR) as compared to the other datasets in this study, making this dataset representative of real-world data. This work only utilises the front six channels, as the back channel was found to be contaminated with breathing noise.
@@ 2.1.3 Generative and augmentation datasets
In addition to training-a, the Icentia dataset [29] was used to provide novel ECG inputs for generating PCG signals. This dataset contains 11,000 patients and 2.77 billion labelled heartbeats sampled at 250 Hz, with 541,794 segments. Each beat is labelled as normal, premature atrial contraction, premature ventricular contraction, or one of several rhythm types (sinus, atrial fibrillation, atrial flutter).
To enhance robustness to noise, augmentations were performed using additional datasets: EPHNOGRAM (for PCG) and the MIT-BIH Noise Stress Test Database (for ECG). EPHNOGRAM comprises PCG recordings from 24 healthy adults during rest and stress conditions [30]. The MIT-BIH dataset contains 12 half-hour ECG recordings and three half-hour recordings of typical noise, including baseline wander, muscle artefacts, and electrode motion [31]. These noise samples were used to augment ECG signals.
@@ 3.1.1 Single channel augmentations
The PCG and ECG augmentation process is shown in Figure 8 (adapted from [32]). Each augmentation has a specific probability of occurring per sample: harmonic-percussive source separation (75%), white noise (7.5%), time stretching (25%), amplitude modulation (75%), baseline wander (75%), parametric equalisation (25%), and clinical noise (50%).
@@ 3.1.2 Multichannel augmentations
The mPCG followed similar augmentations to the single channel augmentations [32], but was further extended from [32] to support mPCG signals, by synchronising the time-stretching across channels the same way as it was for synchronising the PCG and ECG data. All other augmentations were applied independently to each channel. Figure 9 shows the procedure for multichannel PCG data augmentation, with the time-stretching augmentation being synchronised across all channels in the augmented patient.
@@ 3.2.1 Single channel PCG
Synthetic PCG signals were generated using ECG signals from the Icentia dataset [29] as conditioners, following [32]. WaveGrad [8] and DiffWave [7] generated 3,200 patients’ data each, using a 3:1 ratio of normal to abnormal samples. Local conditioning was done using ECG mel-spectrograms, and global conditioning was done using disease labels. The generation process is summarised in Figure 10.
@@
Training lasted 24 hours on an RTX 3090, with cardiac cycle rearrangement (25% probability) applied during training to reduce overfitting. Rearrangement types include shuffling large groups, 1–4 heart cycle chunks, and individual cycles [32]. Crossfading was used to minimise artifacts.
@@ 3.4 Models
From training, the chosen model is based on the model’s performance on the validation set, with the highest Matthew’s correlation coefficient (MCC) value. This value is used as it combines much information about how well the model performs into a single value [33]. While not fully comprehensive, it resulted in the best-performing models when used as the main selection metric. The MCC, as well as other metrics, are shown below and is used to evaluate model performance against the test set, as the MCC value does not fully summarise the confusion matrix. TP represents true positives, TN true negatives, FP false positives, and FN false negatives.
@@
Hyperparameters were tuned using a Bayesian optimisation (Optuna [34]) on the augmented datasets. The average (across five runs, to account for the variation of training the models) MCC score on the validation set was utilised as the optimisation metric. The initial hyperparameters for the baseline model without the augmented dataset are found in Table 4. This optimisation used 10 epochs from the original data with at least 30 augmentations per patient for the single channel and multimodal models, following the first step of the training schedule that the models undergo.

### v3

# Scaling to Multimodal and Multichannel Heart Sound Classification with Synthetic and Augmented Biosignals

@@ Abstract
Cardiovascular diseases (CVDs) are the leading cause of death worldwide, accounting for approximately 17.9 million deaths each year. Early detection is critical, creating a demand for accurate and inexpensive pre-screening methods. Deep learning has recently been applied to classify abnormal heart sounds indicative of CVDs using synchronised phonocardiogram (PCG) and electrocardiogram (ECG) signals, as well as multichannel PCG (mPCG). However, state-of-the-art architectures remain underutilised due to the limited availability of synchronised and multichannel datasets. Augmented datasets and pre-trained models provide a pathway to overcome these limitations, enabling transformer-based architectures to be trained effectively on smaller datasets.
This work combines traditional signal processing with denoising diffusion models, WaveGrad and DiffWave, to create an augmented dataset to fine-tune a Wav2Vec 2.0-based classifier on multimodal and multichannel heart sound datasets.
The approach achieves state-of-the-art performance. On the Computing in Cardiology (CinC) 2016 dataset of single channel PCG, accuracy, unweighted average recall (UAR), sensitivity, specificity and Matthew’s correlation coefficient (MCC) reach 92.48%, 93.05%, 93.63%, 92.48%, and 0.8283, respectively. Using the synchronised PCG and ECG signals of the training-a dataset from CinC, 93.14%, 92.21%, 94.35%, 90.10%, and 0.8380 are achieved for accuracy, UAR, sensitivity, specificity and MCC, respectively. Using a wearable vest dataset consisting of mPCG data, the model achieves 77.13% accuracy, 74.25% UAR, 86.47% sensitivity, 62.04% specificity, and 0.5082 MCC.
These results demonstrate the effectiveness of transformer-based models for CVD detection when supported by augmented datasets, highlighting their potential to advance multimodal and multichannel heart sound classification.
@@ 1 Introduction
Cardiovascular diseases (CVDs) are the leading cause of death globally, accounting for approximately 17.9 million deaths each year [who]. Since CVD treatment is most effective when the condition is detected early, there is a pressing need for accurate and affordable pre-screening methods. Cardiac auscultation is one such technique: it is inexpensive, non-invasive, and widely used, relying on physicians listening to heart sounds to detect abnormalities indicative of CVD. However, auscultation yields relatively low diagnostic accuracy, partly because heart sounds often lie near the threshold of human hearing [heart1, heart2, heart3].
Recent advances in computer-aided methods have demonstrated promise in improving the accuracy of abnormal heart sound classification. Combining phonocardiogram (PCG) and electrocardiogram (ECG) signals within deep learning frameworks has produced favourable results [milan]. However, current datasets suffer from limitations, including class imbalance, low signal-to-noise ratio, and limited size, all hindering robust and accurate classification. Current state-of-the-art (SOTA) approaches often use convolutional neural networks (CNNs), many of which leverage pre-trained image-based architectures with spectrogram inputs. Transformer-based models remain underexplored in this domain, with only Vision Transformers (ViTs) evaluated on spectrogram inputs. Therefore, there is significant potential in applying transformers directly to raw audio signals.
Using raw signals as input offers the advantage of preserving phase information and enabling models to learn features that may be difficult to extract from time-frequency representations. However, learning time-frequency features from raw audio alone can be challenging. Modern architectures such as transformers generally require substantial data to outperform existing approaches [attention]. Transformer-based models that operate on raw signals show potential for improved performance, but their data requirements must be addressed.
Recent advances in synthetic audio generation using diffusion models offer a promising solution to this data scarcity, enabling easier training of data-hungry models and potentially yielding SOTA performance [diffwave, wavegrad]. Traditional data augmentation techniques have also been applied to mitigate issues with limited and imbalanced datasets [augmentation]. Furthermore, fine-tuning large pre-trained models provides another avenue for overcoming data limitations [milan, goutham]. Models like Wav2Vec 2.0 (Wav2Vec2) [wav2vec] have shown excellent performance in various speech classification tasks [dysarthria], making them strong candidates for transfer learning in PCG and ECG classification.
A recent development includes a wearable device equipped with up to seven PCG sensors, enabling the collection of synchronised multichannel PCG data [RongFynnNordholm2023PreScreeningCAD]. This advancement makes it feasible to train models utilising multimodal and multichannel data for abnormal heart sound detection.
@@ 1.1 Phonocardiogram and Electrocardiogram Signals
PCG signals consist of multiple sounds created by sudden changes in blood flow within the heart, causing vibrations [leatham]. The fundamental heart sounds are S1 and S2. S1 occurs due to isovolumetric ventricular contraction at the beginning of systole, and S2 results from the closure of the aortic and pulmonic valves at the beginning of diastole.
While S1 and S2 are the most audible sounds, other sounds can also be heard, such as the third heart sound (S3), fourth heart sound (S4), systolic ejection click, mid-systolic click, opening snap, and heart murmurs, which result from turbulent, fast-flowing blood [cinc]. These sounds all lie in the low-frequency range: S1 ranges from 10–140Hz with the highest energy around 25–45Hz; S2 spans 10–200Hz with energy concentrated around 55–75Hz; and S3 and S4 fall between 20–70Hz, though they are less audible. Murmurs, which may indicate CVDs, can be found in a wider frequency range from 25Hz to 400Hz [schmidt], with some extending up to 600Hz but with less energy [springer].
ECG signals represent the electrical activity of the heart [heart4]. An ECG signal comprises P waves, QRS complexes, and T waves, with a U wave occasionally present [heart5]. These components contain diagnostic information useful in identifying CVDs. ECG signals are typically filtered between 0.5Hz and 40Hz to remove baseline wander and unwanted noise [ecgfiltering]. In patients with coronary artery disease (CAD), symptoms such as T-wave inversion, ST-T abnormalities, left ventricular hypertrophy, and premature ventricular contractions have been documented [ecgcad].
Combining PCG and ECG signals has yielded better results than using either signal alone [milan], as both contain complementary diagnostic features. Figure 1 illustrates these signals in a patient with mitral valve prolapse.
@@
Combining multiple PCG sensors in different auscultation regions can yield better results than single-channel PCG, as each channel offers a different resolution of the region where the murmur originates [vestpaper]. Figure 2 shows a CAD patient with multichannel PCG (mPCG) data.
@@
Wav2Vec 2.0 (Wav2Vec2), developed by Meta, is a speech-to-text model with a transformer-based architecture. It consists of three primary components: a feature encoder, transformer encoder and a quantisation module for pre-training [wav2vec].
@@
The quantisation module discretises and then encodes the output using a Gumbel-Softmax function during self-supervised training. Training proceeds in two phases: first, self-supervised training on large unlabelled audio datasets (e.g., LibriSpeech [librispeech]) to learn speech representations; second, fine-tuning with 960 hours of labelled speech data across diverse accents and languages [wav2vec].
@@ 1.3 Diffusion Models
Diffusion probabilistic models are generative models that convert a simple distribution (e.g., isotropic Gaussian) into a complex data distribution via a Markov chain process [sohldickstein, ho2020]. Training is based on optimising the variational lower bound (ELBO), even when the data likelihood is intractable.
These models, applied in audio and image synthesis, relate to denoising score matching and do not require separate encoder or discriminator networks like VAEs [kingma2014] or GANs [goodfellow2014]. This advantage avoids issues such as posterior collapse and mode collapse, making diffusion models particularly effective for high-fidelity audio generation by whitening training data latents with a parameter-free noise process.
@@ 1.3.1 WaveGrad
WaveGrad is a diffusion model for conditional audio synthesis [wavegrad]. It employs upsampling (UBlocks) and downsampling (DBlocks) blocks, conditioned on mel-spectrogram inputs, along with Feature-wise linear modulation (FiLM) modules. These blocks resemble those in the GAN-TTS model [gan-tts]. Figure 4 shows the overall architecture. Module components are shown in Figure 5. The loss function measures the difference between the added noise in the forward diffusion and the predicted noise during denoising.
@@ 1.3.2 DiffWave
DiffWave is a diffusion model for raw audio synthesis with both conditional and unconditional variants [diffwave]. The model uses 1D convolutions and fully connected layers, with its core comprising bi-directional dilated convolutions and residual connections. It is trained using a single ELBO-based objective without auxiliary losses. Conditional generation uses local conditioning signals and global conditioning via discrete labels.
@@ 2.1.1 Multimodal dataset
The classification dataset is sourced from the 2016 PhysioNet Computing in Cardiology Challenge (CinC), comprising five databases (training-a to training-e) [cinc]. There are 3153 recordings sourced from 764 patients [cinc], some of which are very noisy. Of these recordings, 665 are abnormal and 2488 are normal, with each recording lasting 5–120 seconds. Database training-a includes synchronised ECG and PCG recordings; out of 409 recordings, 405 contain both signal types (288 abnormal, 117 normal, hence, we conduct analysis on PCG signal only models using the entire dataset, and PCG-ECG hybrid models using only training-a.
Table 1: Datasets from CinC 2016 (adapted from [cinc])
@@ 2.1.2 Multichannel dataset
Recordings were obtained from subjects using a multichannel wearable vest with seven phonocardiogram sensors recording at different auscultation sites [vestpaper]. Ninety-six subjects were diagnosed with coronary artery disease through angiography, and an additional 61 were subjects without the disease. Of the 61, 21 were control subjects below the age of 35, assumed to be normal, as the risk of CAD is significantly higher in people aged 45 and over. The multichannel device had sixty seconds of data recorded in a hospital with background noise, with normal breathing occuring. This vest dataset also did not have optimal positioning of all the stethoscopes, resulting in lower signal-to-noise ratio (SNR) as compared to the other datasets in this study, making this dataset representative of real-world data. This work only utilises the front six channels, as the back channel was found to be contaminated with breathing noise.
@@ 2.1.3 Generative and augmentation datasets
In addition to training-a, the Icentia dataset [icentia] was used to provide novel ECG inputs for generating PCG signals. This dataset contains 11,000 patients and 2.77 billion labelled heartbeats sampled at 250 Hz, with 541,794 segments. Each beat is labelled as normal, premature atrial contraction, premature ventricular contraction, or one of several rhythm types (sinus, atrial fibrillation, atrial flutter).
To enhance robustness to noise, augmentations were performed using additional datasets: EPHNOGRAM (for PCG) and the MIT-BIH Noise Stress Test Database (for ECG). EPHNOGRAM comprises PCG recordings from 24 healthy adults during rest and stress conditions [ephnogram]. The MIT-BIH dataset contains 12 half-hour ECG recordings and three half-hour recordings of typical noise, including baseline wander, muscle artefacts, and electrode motion [mit]. These noise samples were used to augment ECG signals.
@@ 3.1.1 Single channel augmentations
The PCG and ECG augmentation process is shown in Figure 8 (adapted from [epcgdiffusion]). Each augmentation has a specific probability of occurring per sample: harmonic-percussive source separation (75%), white noise (7.5%), time stretching (25%), amplitude modulation (75%), baseline wander (75%), parametric equalisation (25%), and clinical noise (50%).
@@ 3.1.2 Multichannel augmentations
The mPCG followed similar augmentations to the single channel augmentations [epcgdiffusion], but was further extended from [epcgdiffusion] to support mPCG signals, by synchronising the time-stretching across channels the same way as it was for synchronising the PCG and ECG data. All other augmentations were applied independently to each channel. Figure 9 shows the procedure for multichannel PCG data augmentation, with the time-stretching augmentation being synchronised across all channels in the augmented patient.
@@ 3.2.1 Single channel PCG
Synthetic PCG signals were generated using ECG signals from the Icentia dataset [icentia] as conditioners, following [epcgdiffusion]. WaveGrad [wavegrad] and DiffWave [diffwave] generated 3,200 patients’ data each, using a 3:1 ratio of normal to abnormal samples. Local conditioning was done using ECG mel-spectrograms, and global conditioning was done using disease labels. The generation process is summarised in Figure 10.
@@
Training lasted 24 hours on an RTX 3090, with cardiac cycle rearrangement (25% probability) applied during training to reduce overfitting. Rearrangement types include shuffling large groups, 1–4 heart cycle chunks, and individual cycles [epcgdiffusion]. Crossfading was used to minimise artifacts.
@@ 3.4 Models
From training, the chosen model is based on the model’s performance on the validation set, with the highest Matthew’s correlation coefficient (MCC) value. This value is used as it combines much information about how well the model performs into a single value [mcc]. While not fully comprehensive, it resulted in the best-performing models when used as the main selection metric. The MCC, as well as other metrics, are shown below and is used to evaluate model performance against the test set, as the MCC value does not fully summarise the confusion matrix. TP represents true positives, TN true negatives, FP false positives, and FN false negatives.
@@
Hyperparameters were tuned using a Bayesian optimisation (Optuna [optuna]) on the augmented datasets. The average (across five runs, to account for the variation of training the models) MCC score on the validation set was utilised as the optimisation metric. The initial hyperparameters for the baseline model without the augmented dataset are found in Table 4. This optimisation used 10 epochs from the original data with at least 30 augmentations per patient for the single channel and multimodal models, following the first step of the training schedule that the models undergo.